OpenStudy (anonymous):

Guys, I need help on this one: \int((5x^3-3x*2+2x-1)/(x^4-x^2))

OpenStudy (luigi0210):

Is it this: \[\int\limits \frac{5x^3-3x^2+2x-1}{x^4-x^2}\]

OpenStudy (tkhunny):

When was the last time you warmed up your Partial Fractions?

OpenStudy (anonymous):

@tkhunny I did yesterday, but still can't find the solution....

OpenStudy (tkhunny):

You'll need some denominators. Looks like x, x^2, x+1 and x-1. Now what?

OpenStudy (anonymous):

I'll retry that, please wait a moment ;)

OpenStudy (anonymous):

@tkhunny : Still can't find the solution: \[\frac Ax+\frac{B}{x^2}+\frac{C}{x+1}+\frac{D}{x-1}=Ax(x+1)(x-1)+B(x+1)(x-1)+Cx^2(x-1)+Dx^2(x+1) \] I found illogical solution for B, B=-1 and B=1 :? Can you help please?

OpenStudy (tkhunny):

Did you continue to expand your expression? There aren't too many shortcuts, here.

OpenStudy (anonymous):

"Did you continue to expand your expression?" I did expand the RHS above and equate them with 5x^3-3x*2+2x-1, that way I found B=-3 and B=1 too. Do you mean that?

OpenStudy (tkhunny):

You should get: \(x^{3}(A+C+D) + x^{2}(B+C-D) + x(-A) + (-B)\) Did you?

OpenStudy (anonymous):

I thought I messed up somewhere, I got \[(A+C+D)x^3+Bx^2+(-A-C+D)x-B\]

OpenStudy (tkhunny):

Better try that again. That would indeed lead to an irrational solution. Be careful. Make an accounting system so that you are not confused.

OpenStudy (anonymous):

Can you lead me step by step in expanding the fraction? I really don't know where I messed up :?

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

This is a calculus problem. You should not be hung up on the algebra, tedious though it may be. See how I wrote the variable in front of the coefficient? This is how I collect coefficients from annoyingly-long expressions. Develop some sort of scheme that benefits you. Keep trying until you get it!

OpenStudy (anonymous):

Oh myy..... I just slupped \[Cx^2(x-1)=Cx^3-Cx\] Thanks mate, here's your medal ;)

OpenStudy (anonymous):

*slipped

OpenStudy (tkhunny):

So, what did you get? A = ? B = ? C = ? D = ? There are lots more errors to make, and we haven't even gotten to the calculus.

OpenStudy (anonymous):

Lol, okay, I'll try again:

OpenStudy (tkhunny):

\(A(x^{3} - x)\)

OpenStudy (tkhunny):

\(B(x^{2} - 1)\)

OpenStudy (tkhunny):

\(C(x^{3} + x^{2})\)

OpenStudy (tkhunny):

\(D(x^{3} - x^{2})\)

OpenStudy (tkhunny):

Start collecting!

OpenStudy (anonymous):

A=-2, B=1, C=3/2, D=11/2. Is that it?

OpenStudy (tkhunny):

I don't know that you didn't just copy that out of a book. Shall we ignore that and get to the calculus?

OpenStudy (anonymous):

Lol, okay, we get to calculus. I think I can do the A, C, and D part with substitution? how about the B part: \[\int{\frac{1}{x^2}}dx\]

OpenStudy (tkhunny):

There's no substitution. Do this one first \(\int \dfrac{1}{x}\;dx\)

OpenStudy (anonymous):

It's ln(x) +C right?

OpenStudy (tkhunny):

Not quite. That's why I asked. Without any clue about x, the correct result is \(\int \dfrac{1}{x}\;dx = ln|x| + C\) The absolute values are not optional.

OpenStudy (anonymous):

Ooh, any reason for the absolute mark?

OpenStudy (anonymous):

Then for C and D, \[\frac 32 \ln{|x+1|}+\frac{11}{2}\ln|x-1|+C\] amI right?

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

If you KNOW x > 0, you can skip the absolute values. If you do NOT know x > 0, they are mandatory because the logarithm function blows up!

OpenStudy (tkhunny):

B is just the power rule or polynomial rule. Whatever you wish to call it.

OpenStudy (anonymous):

Okay, thanks mate. You saved my day :)

OpenStudy (tkhunny):

You may also use your logarithm properties to simplify the expression.